Find the value of y for a given value of x, if y varies dirctly with x
If y = -252 when x = 84, what is y when x = 74?

Answer by stanbon(75887)

Find the value of y for a given value of x,
if y varies dirctly with x
y = k*x
-----------
If y = -252 when x = 84
-252 = k*84
k = -3
--------
Equation:
y = -3x
------
What is y when x = 74?
y = -3*74 = -222
==================
Cheers,
Stan H.
==================